OpenStudy (anonymous):

Find the slope given the points (4, -6) and (-13, 5) I got m= -11/17

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

thanks.

hero (hero):

You're like the first student I've seen that actually posts your possible answer

OpenStudy (anonymous):

wtf, arent you asian?

hero (hero):

Asian American maybe

OpenStudy (anonymous):

asians are geniuses

hero (hero):

She's just checking her answers.... At least she's posting her possible answers.... Normally students don't do that so it's commendable

OpenStudy (anonymous):

slope =11/17

hero (hero):

I get slope is \[-\frac{11}{17}\]

OpenStudy (anonymous):

i got the same answer good job

OpenStudy (anonymous):

Find the slope given the points (4, -6) and (-13, 5) 5-(-6)/(-13-4)=11/(-17) = -11/17 correct